A is a roller and C is pin connected
3-9. Determine the force in each member of the truss. State if the members are in tension or compression. E 12 kN 8 kN 3 m 45° 24 kN 4 m Prob. 3-9
Answer #1

1-9 ame uon Rc A Sins Rc 19.KN FAG 3.8y-o 12.5=FAF sine Sim36-83 FA 20-83 Kn (Cnp) 12 FA-FAF FAo 19-20-33 36 FefFer FDE AP, roc ou KN FEF FED = 16-67.-19. 18C63637 FE D = 1.G否KNr.muen FEC 198 sio 36-37 1-66 sin 36 2vN FLD